1-2 parallel very long wires 1 and 2 are separated by a distance of 0,065m and carry currents I1 of 15A and I2=7A.What is the magnitude of the force that the magnetic field that wire 1 applies to a 1,5m of wire 2 when the currents are in opposite direction.
2-A 150 turn solenoid carries a current of 12A,the radius of the solenoid is 0,050m and its strength is 0,18m.what is the magnetic flux through the circular crodd-sectional area at the center of the solenoid?
1
Expert's answer
2020-04-29T09:46:02-0400
  1. As per the given question,

There is two parallel wires,

distance between the wires (d)=0.065m

Current in the wire 1, "(i_1)=15A"

Current in the wire 2, "(i_2)=7A"

Length of the wire "(l)=1.5m"

Now,

"F=\dfrac{\mu_o i_1}{4\pi d}i_2 l"


"\Rightarrow F=10^{-7}\times 15\times 7\times 1.5\times \dfrac{1}{0.065}=2.432\times 10^{-4}N"

2.As per the data given in the question,

Number of turns in the solenoid "(N)=150"

Current in the wires (i)=12A

Radius of the solenoid (r)=0.05m

Length of the solenoid (l)=0.18m

magnetic field strength per turn "H=150\times \dfrac{12}{0.18}T=1\times 10^{4}T"


"\Rightarrow B=\mu H=4\pi \times 10^{-7}\times 1\times 10^{4}T=4\pi \times 10^{-3}T"

Hence, magnetic flux

"\Rightarrow \phi=BA=4\pi \times 10^{-3}\times \pi \times (0.05)^2=2.5\times 10^{-3}\times4\pi \times 10^{-3}Tm^2"

"\Rightarrow \phi =31.41\times 10^{-6}Tm^2"